OpenStudy (anonymous):

reduce the equation 3x+y-2=0 to the slope-ntercept form

OpenStudy (joannablackwelder):

Slop intercept form is y=mx + b

OpenStudy (joannablackwelder):

So we need to get y by itself on one side of the equals sign and everything else on the other side.

OpenStudy (joannablackwelder):

Any ideas @velasco ?

OpenStudy (anonymous):

yes.

OpenStudy (joannablackwelder):

What do you think?

OpenStudy (anonymous):

you can add equation,attachments,and drawings to better explain your answer.

OpenStudy (joannablackwelder):

I can't because I'm on an iPad. Did you have any thoughts on how to solve for y?

OpenStudy (anonymous):

it is correct that solve like this,y=3x+2?

OpenStudy (joannablackwelder):

Very close" to move the 3x you need to subtract on both sides.

OpenStudy (joannablackwelder):

y = -3x + 2

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (anonymous):

thank you....

OpenStudy (joannablackwelder):

You're welcome.

OpenStudy (joannablackwelder):

Welcome to OpenStudy!

OpenStudy (anonymous):

so how to solve like this one?2x-7y+5=0?

OpenStudy (joannablackwelder):

Similarly, but an extra step since y has a coefficient.

OpenStudy (joannablackwelder):

Can you first solve for -7y?

OpenStudy (anonymous):

so like this?7y=2x+5 or y=2/7x+5/7

OpenStudy (joannablackwelder):

Yes, good!

OpenStudy (joannablackwelder):

You catch on fast! :)

OpenStudy (anonymous):

thank you again......

OpenStudy (joannablackwelder):

No worries. :)